OpenStudy (anonymous):

partial fractions.. x/x^2-4

OpenStudy (anonymous):

Write it as \[ \frac{x}{x^2-4}=\frac{A}{x-2}+\frac{B}{x+2} \]

OpenStudy (anonymous):

Find A and B

OpenStudy (anonymous):

Multiply the two sides by (x-2) to get \[ \frac{x (x-2)}{(x-2) (x+2)}=\frac{A (x-2)}{x-2}+\frac{B (x-2)}{x+2}\\ \frac{x}{x+2}=\frac{A}{1}+\frac{B (x-2)}{x+2} \] Put x =2 and get A Do the same for B

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

The final answer looks like \[ \frac{x}{x^2-4}=\frac{1}{2 (x+2)}+\frac{1}{2 (x-2)} \]

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

thats the final answer

OpenStudy (anonymous):

yes